OpenStudy (anonymous):

Let f(x) = x2 - 7x + 3 and g(x) = x + 4. Find f(x) + g(x).?

OpenStudy (anonymous):

x2-6x+7

OpenStudy (anonymous):

f(x)+g(x)=(x^2-7x+3)+(x+4) =x^2-6x+7